# Self-consistent inclusion of disorder in the BCS-BEC crossover near $T_{c}$

### v1

@@ Abstract
We develop a systematic theoretical approach to incorporate the effects of a static white‐noise disorder into the BCS-BEC crossover near the critical temperature ($T_{c}$) of the superfluid transition. Starting from a functional‐integral formulation in momentum-frequency space, we derive an effective thermodynamic potential that fully accounts for Gaussian fluctuations of the order‐parameter field and its coupling to the disorder potential. The effective action, expanded to second order in both the disorder potential and the bosonic field, naturally involves third- and fourth-order terms arising from the logarithmic expansion near $T_{c}$. This formalism, valid across the entire BCS-BEC crossover, reproduces the well-established BCS and BEC limits and yields self-energy expressions consistent with previous analyses for non-interacting point bosons and tightly-bound fermion pairs. The approach applies equally to continuum and lattice systems and provides a natural framework for generalizations to multiband models.
@@ I Introduction
The study of disordered quantum gases has emerged as a powerful setting for exploring the interplay among interactions, coherence and localization in highly controllable environments  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Ultracold Fermi gases, in particular, provide an ideal platform for realizing the BCS-BEC crossover under tunable conditions [13, 14, 15, 16], where both the interaction strength and the disorder potential can be precisely controlled via Feshbach resonances and optical speckle fields. While the crossover is well characterized in clean systems, the combined effects of disorder and strong pairing correlations remain only partially understood [4, 7, 10, 11]. Theoretically, Fermi gases interpolate between two qualitatively distinct limits: on the BCS side, Anderson’s theorem guarantees the robustness of $T_{c}$ against weak nonmagnetic disorder, whereas on the BEC side, the transition is governed by phase coherence among tightly-bound fermion pairs, which is highly susceptible to localization and incoherence [17, 18, 19, 20]. A unified finite-temperature description valid throughout the BCS-BEC crossover is therefore crucial for interpreting cold-atom experiments and for connecting continuum and lattice models of disordered superfluidity. Previous theoretical studies have either focused on the zero-temperature ($T=0$) superfluid phase [17, 20] or determined $T_{c}$ through semiclassical analyses performed within the local-density approximation and the replica trick [18]. Other studies have employed diagrammatic methods at the lowest nontrivial order for the normal phase above $T_{c}$, complemented by accurate numerical evaluations of the corresponding diagrammatic contributions [19]. A fully self-consistent treatment valid across the entire crossover region at finite temperatures is still lacking.
@@
Building on a functional-integral framework in momentum-frequency space, we construct an effective thermodynamic potential that incorporates Gaussian fluctuations of the order parameter together with their coupling to a static disorder potential. Crucially, our formulation provides a fully quantum and systematically controlled approach that captures the effects of weak disorder across the BCS-BEC crossover in the vicinity of $T_{c}$, without resorting to the local-density approximation or the replica trick. By expanding the effective action to second order in both the disorder potential and the bosonic field, and by retaining the necessary third- and fourth-order terms in the logarithmic expansion, our approach yields a controlled finite-temperature description that bridges the BCS and BEC regimes. In contrast to the $T=0$ approach, where disorder couples to bosonic fluctuations via the second-order term [17, 20], these contributions vanish near $T_{c}$, making higher-order terms essential. The formalism reproduces known limiting behaviors for weakly-interacting fermions and bosons, recovers the fermionic and bosonic self-energy diagrams of Ref. [19] for the three-dimensional continuum model, and extends beyond them. Applicable to both continuum and lattice systems, the theory provides a natural route for describing the loss of phase coherence near $T_{c}$ and for connecting microscopic disorder models in ultracold gases with those relevant to superconductors.
@@ II.2 Effective Gaussian action
which is a formally exact expression, where the trace $\mathrm{Tr}$ is taken over both Nambu and momentum-frequency spaces. Since the saddle-point order parameter vanishes (i.e., $\Delta\to 0$) near $T_{c}$, the inverse Nambu-Gorkov Green’s function can be written as $\mathbf{G}^{-1}_{kk^{\prime}}=\boldsymbol{\mathcal{G}}^{-1}_{k}\delta_{kk^{\prime}}-\boldsymbol{\Sigma}_{kk^{\prime}},$ where
@@
and remains unaffected by disorder [17, 18, 20]. Since the disorder correction to the second-order action, $\mathcal{S}_{2}^{d}=\frac{1}{2\beta N}\sum_{kq}\mathrm{tr}\big(\boldsymbol{\mathcal{G}}_{k}\boldsymbol{\tau}_{z}\boldsymbol{\mathcal{G}}_{k+q}\boldsymbol{\tau}_{z}\big)V_{-q}V_{q},$ does not couple to $\phi_{q}$, the effective thermodynamic potential depends only on its disorder average, $\langle\mathcal{S}_{2}^{d}\rangle_{d}=\frac{\kappa}{2N}\sum_{kq,q_{\ell}=0}(\mathcal{G}_{k}\mathcal{G}_{k+q}+\mathcal{G}_{-k}\mathcal{G}_{-k-q}).$ In the third-order action
@@ II.3 Effective thermodynamic potential
Although its form appears quite different, the physical origin of this term is the same as that of the corresponding $T=0$ result [17, 20].
@@
where $\Omega_{\mathrm{B}}^{d,1}=\Omega_{\mathrm{B}}^{d,1a}+\Omega_{\mathrm{B}}^{d,1b}.$ Unlike the analogous disorder correction at $T=0$ [17, 20], which has a remarkably simple form, these expressions are considerably more involved because they arise from a higher-order logarithmic expansion. In Sec. IV, we show that while $\Omega_{\mathrm{F}}^{d}$ provides the leading-order disorder correction in the BCS limit, it is $\Omega_{\mathrm{B}}^{d,2}$ that dominates in the BEC limit.
@@ IV.3 Number equation in the BEC limit
which corresponds to twice the standard number equation for a non-interacting spin-0 Bose gas, i.e., $n_{\mathrm{B}}^{0}=2n_{M}^{0}$. Similarly, from Eqs. (20) - (22), we obtain the disorder corrections to the number equation in the BEC limit, noting that the dominant contribution originates from the term containing the largest number of $\Gamma_{q}$ factors. For instance, in Eq. (22) for $\Omega_{\mathrm{B}}^{d,2}$, assuming $|\mu|\gg\Lambda^{2}/(2m)$ and $|\mu|\gg|q_{\ell}|$ for the relevant bosonic energy scales [19], we may approximate $\sum_{p}\mathcal{G}_{q+p}\mathcal{G}_{k+q+p}\mathcal{G}_{-p}\to\sum_{p}\mathcal{G}_{p}^{2}\mathcal{G}_{-p},$ and use Eq. (33), which leads to $\Omega_{\mathrm{B}}^{d,2}\to-\frac{2\kappa}{\beta N^{2}}\sum_{qk,k_{n}=0}\frac{\Gamma_{q}\Gamma_{q+k}}{Z^{2}}.$ Thus, we identify
@@
Thus, we conclude that while the leading-order self-energy diagrams associated with our bosonic contributions $\Omega_{\mathrm{B}}^{d,2}$ and $\Omega_{\mathrm{B}}^{d,1a}$ are properly included in Ref. [19], the diagram corresponding to the remaining bosonic term $\Omega_{\mathrm{B}}^{d,1b}$ is absent from their analysis. However, as discussed in Ref. [19], one finds that $\Sigma_{q}^{\mathrm{B}d,1a}\ll\Sigma_{q}^{\mathrm{B}d,2}$ in the BEC limit, and the same hierarchy holds for the remaining contribution $\Sigma_{q}^{\mathrm{B}d,1b}.$ Therefore, the leading-order effects of weak disorder can be incorporated into the number equation through the self-energy as
@@ V Critical temperature $T_{c}$
It turns out that the effect of the disorder potential on $T_{c}$ differs markedly between a weakly-interacting Fermi gas in the BCS limit and a weakly-interacting Bose gas of tightly-bound fermion pairs in the BEC limit [18, 19]. While the self-energy expressions in Eqs. (24) and (26) may appear superficially similar, a key difference emerges: in the fermionic case, the chemical potential $\mu$ is large and positive at $T_{c}$, effectively shielding the system from disorder. As a result, $T_{c}$ is primarily determined by the saddle-point condition Eq. (12), and Anderson’s theorem protects it from weak disorder. In other words, the presence of a large Fermi surface restricts scattering processes, ensuring that weak disorder has only a minimal effect. In contrast, for tightly-bound fermion pairs in the BEC limit, the effective bosonic chemical potential $\mu_{M}\to 0^{+}$ at $T_{c}$, leaving them fully exposed to disorder. Since $T_{c}$ is essentially set by the number equation Eq. (37), rather than by Eq. (12) which fixes $\mu\to-\varepsilon_{b}/2$, weak disorder directly affects phase coherence through the emergence of an incoherent component in the pair propagator, characterized by the $\sqrt{-iq_{\ell}}$ dependence of the bosonic self-energy $\Sigma_{q}^{M}$. This distinction highlights the fundamental role of quantum statistics in determining the robustness of phase coherence against the disorder potential.
@@
On the other hand, in the BEC limit, direct evaluation of Eq. (34) gives $n_{\mathrm{B}}^{0}=2[m_{M}T_{c}/(2\pi)]^{3/2}\zeta(3/2),$ where $\zeta(3/2)=2.612$ is the Riemann-zeta function, and the leading-order disorder correction is given by Eq. (35). The latter can be calculated either by taking the derivative of Eq. (26) with respect to $\mu_{M}$, or equivalently by performing the momentum sum over $\mathbf{q}$ via direct integration, using $\int_{0}^{\infty}\frac{\sqrt{\varepsilon}\,d\varepsilon}{(iq_{\ell}-\varepsilon)^{2}}=\frac{\pi}{2\sqrt{-iq_{\ell}}}$ for $q_{\ell}\neq 0$. This yields $n_{\mathrm{B}}^{d}=\frac{\kappa_{M}m_{M}^{3}}{2\pi^{2}}T_{c},$ where we set $\mu_{M}\to 0$ and employ $\sum_{q_{\ell}\neq 0}e^{iq_{\ell}0^{+}}=-1.$ The latter identity follows from $\frac{1}{\beta}\sum_{q_{\ell}}e^{iq_{\ell}0^{+}}=\frac{1}{2\pi i}\oint e^{z0^{+}}n_{\mathrm{B}}(z)dz=0,$ since the integrand vanishes along the circular periphery when the contour encircling the poles is extended to infinity. Note that the $q_{\ell}=0$ frequency does not contribute to the disorder correction, and the arising $1/\sqrt{-iq_{\ell}}$ factor is exactly canceled by the $\sqrt{-iq_{\ell}}$ dependence of the bosonic self-energy. See Appendix A for further discussion. As a result, we find
@@
In the BCS-BEC crossover problem, one must solve the saddle-point condition in Eq. (12) together with the number equation in Eq. (27) self-consistently to determine $T_{c}$ and $\mu$ as functions of $U$. Even in the clean case, this procedure is already nontrivial in the crossover regime [21]; the inclusion of disorder makes it substantially more involved due to the complicated forms of the corrections in Eqs. (20) - (22). In Ref. [19], a thorough numerical analysis of the effects of resulting leading-order self-energy diagrams was performed under various approximations across the BCS-BEC crossover region, along with analytic estimates in the extreme BCS and BEC limits. In Ref. [19], it was also shown that including disorder effects at the level of Eqs. (30) and (37) is sufficient for a qualitative understanding of the problem in the intermediate regime [18], although quantitative results depend sensitively on the specific procedure used to handle the cutoff-related issues [19].

### v2

# Self-consistent inclusion of disorder in the BCS-BEC crossover near the critical temperature

@@ Abstract
We develop a systematic theoretical approach to incorporate the effects of a static white-noise disorder into the BCS-BEC crossover near the critical temperature ($T_{c}$) of the superfluid transition. Starting from a functional-integral formulation in momentum-frequency space, we derive an effective thermodynamic potential that fully accounts for Gaussian fluctuations of the order-parameter field and its coupling to the disorder potential. The effective action, expanded to second order in both the disorder potential and the bosonic field, naturally involves third- and fourth-order terms arising from the logarithmic expansion near $T_{c}$. By providing a controlled description of pairing fluctuations and disorder effects, this formalism correctly recovers the well-established BCS and BEC limits. This ensures a consistent physical foundation for analyzing the entire BCS-BEC crossover, effectively anchoring the intermediate regime between these two analytically robust endpoints. The approach applies equally to continuum and lattice systems and provides a natural framework for generalizations to multiband models.
@@ I Introduction
The study of disordered quantum gases has emerged as a powerful setting for exploring the interplay among interactions, coherence and localization in highly controllable environments  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Ultracold Fermi gases, in particular, provide an ideal platform for realizing the BCS-BEC crossover under tunable conditions [14, 15, 16, 17], where both the interaction strength and the disorder potential can be precisely controlled via Feshbach resonances and optical speckle fields. While the crossover is well characterized in clean systems, the combined effects of disorder and strong pairing correlations remain only partially understood [4, 11, 12]. Theoretically, Fermi gases interpolate between two qualitatively distinct limits: on the BCS side, Anderson’s theorem guarantees the robustness of $T_{c}$ against weak nonmagnetic disorder, whereas on the BEC side, the transition is governed by phase coherence among tightly-bound fermion pairs, which is highly susceptible to localization and incoherence [18, 19, 20, 21]. A unified finite-temperature description valid throughout the BCS-BEC crossover is therefore crucial for interpreting cold-atom experiments and for connecting continuum and lattice models of disordered superfluidity. Previous theoretical studies have either focused on the zero-temperature ($T=0$) superfluid phase [18, 21] or determined $T_{c}$ through semiclassical analyses performed within the local-density approximation and the replica trick [19]. Other studies have employed diagrammatic methods at the lowest nontrivial order for the normal phase above $T_{c}$, complemented by accurate numerical evaluations of the corresponding diagrammatic contributions [20], or have relied on $T$-matrix approaches [22]. A fully self-consistent treatment valid across the entire crossover region at finite temperatures is still lacking.
@@
Building on a functional-integral framework in momentum-frequency space, we construct an effective thermodynamic potential that incorporates Gaussian fluctuations of the order parameter together with their coupling to a static disorder potential. Crucially, our formulation provides a fully quantum and systematically controlled approach that captures the effects of weak disorder throughout the BCS-BEC crossover at leading order in the vicinity of $T_{c}$, without resorting to the local-density approximation or the replica trick. By expanding the effective action to second order in both the disorder potential and the bosonic field, and by retaining the necessary third- and fourth-order terms in the logarithmic expansion, our approach yields a controlled finite-temperature description that bridges the BCS and BEC regimes. In contrast to the $T=0$ approach, where disorder couples to bosonic fluctuations via the second-order term [18, 21], these contributions vanish near $T_{c}$, making higher-order terms essential. The formalism reproduces known limiting behaviors for weakly-interacting fermions and bosons, recovers the fermionic and bosonic self-energy diagrams of Ref. [20] for the three-dimensional continuum model, and extends beyond them. Applicable to both continuum and lattice systems, the theory provides a natural route for describing the loss of phase coherence near $T_{c}$ and for connecting microscopic disorder models in ultracold gases with those relevant to superconductors.
Experimental studies of disordered ultracold Fermi gases in the regime of static disorder across the BCS-BEC crossover remain relatively limited [11, 12]. Nevertheless, measurements of dipole oscillations in static speckle potentials have demonstrated that the damping coefficient depends sensitively on the interaction strength, indicating a nontrivial interplay between disorder and pairing correlations [11]. More recently, experiments have explored the interaction-dependent stability of superfluid coherence under controlled optical speckle disorder, including studies of quenched-disorder dynamics [12]. These works highlight the strong sensitivity of interacting superfluids to disorder and demonstrate the feasibility of probing disorder effects across different interaction regimes. The present theory addresses the complementary regime of weak static disorder near $T_{c}$. In this limit, we predict a qualitative change in the disorder-induced shift of $T_{c}$ across the BCS-BEC crossover: a slight enhancement of $T_{c}$ on the BCS side and a suppression on the BEC side. This sign change provides a clear equilibrium benchmark for future experiments capable of measuring $T_{c}$ shifts under controlled weak disorder.
@@ II.2 Effective Gaussian action
which is a formally exact expression, where the trace $\mathrm{Tr}$ is taken over both Nambu and momentum-frequency spaces. Since the saddle-point order parameter vanishes (i.e., $\Delta\to 0$) near $T_{c}$, the inverse Nambu-Gor’kov Green’s function can be written as $\mathbf{G}^{-1}_{kk^{\prime}}=\boldsymbol{\mathcal{G}}^{-1}_{k}\delta_{kk^{\prime}}-\boldsymbol{\Sigma}_{kk^{\prime}},$ where
@@
and remains unaffected by disorder [18, 19, 21]. Since the disorder correction to the second-order action, $\mathcal{S}_{2}^{d}=\frac{1}{2\beta N}\sum_{kq}\mathrm{tr}\big(\boldsymbol{\mathcal{G}}_{k}\boldsymbol{\tau}_{z}\boldsymbol{\mathcal{G}}_{k+q}\boldsymbol{\tau}_{z}\big)V_{-q}V_{q},$ does not couple to $\phi_{q}$, the effective thermodynamic potential depends only on its disorder average, $\langle\mathcal{S}_{2}^{d}\rangle_{d}=\frac{\kappa}{2N}\sum_{kq,q_{\ell}=0}(\mathcal{G}_{k}\mathcal{G}_{k+q}+\mathcal{G}_{-k}\mathcal{G}_{-k-q}).$ Here $\boldsymbol{\tau}_{z}$ denotes the $z$ component of the Pauli matrices acting in the Nambu-Gor’kov space. In the third-order action
@@ II.3 Effective thermodynamic potential
Although its form appears quite different, the physical origin of this term is the same as that of the corresponding $T=0$ result [18, 21]. It corresponds to the Feynman diagram shown in Fig. 1(a).
Figure 1: The diagrammatic representations of the disorder-induced thermodynamic potentials, given by Eqs. (18), (20), (21), and (22), offer an intuitive interpretation of the main theoretical results. In these diagrams, solid, dashed, and double lines represent $\mathcal{G}_{k}$, $V_{q}$, and $\Gamma_{q}$, respectively.
@@
where $\Omega_{\mathrm{B}}^{d,1}=\Omega_{\mathrm{B}}^{d,1a}+\Omega_{\mathrm{B}}^{d,1b}.$ Unlike the analogous disorder correction at $T=0$ [18, 21], which has a remarkably simple form, these expressions are considerably more involved because they arise from a higher-order logarithmic expansion. They correspond to the Feynman diagrams shown in Figs. 1(b)-1(d), which admit a clear and intuitive diagrammatic interpretation   [^1] We are grateful to the anonymous referee for pointing out that each of our disorder contributions can be naturally associated with such intuitive Feynman diagrams . In Sec. IV, we show that while $\Omega_{\mathrm{F}}^{d}$ provides the leading-order disorder correction in the BCS limit, it is $\Omega_{\mathrm{B}}^{d,2}$ that dominates in the BEC limit.
@@ IV.1 Number equation in the BCS limit
It is important to clarify the scope of this analysis. In Eq. (30), the fermionic self-energy is retained only through its real, momentum-independent contribution, which yields an ultraviolet-divergent shift of the chemical potential. The imaginary part of the self-energy, associated with quasiparticle lifetime effects induced by disorder [20], is neglected within this leading-order near-$T_{c}$ treatment. While such lifetime effects and the associated vertex corrections in the particle-particle ladder (Cooperon) channel play a central role in microscopic diagrammatic formulations of Anderson’s theorem for disordered BCS superconductors, their self-consistent inclusion lies beyond the Gaussian fluctuation framework employed here. Consequently, the present results should be understood as describing the robustness of the number equation and the transition temperature against weak disorder near $T_{c}$, rather than as a full treatment of quasiparticle lifetimes or transport properties in the weak-coupling BCS regime.
@@ IV.3 Number equation in the BEC limit
which corresponds to twice the standard number equation for a non-interacting spin-0 Bose gas, i.e., $n_{\mathrm{B}}^{0}=2n_{M}^{0}$. Similarly, from Eqs. (20) - (22), we obtain the disorder corrections to the number equation in the BEC limit, noting that the dominant contribution originates from the term containing the largest number of $\Gamma_{q}$ factors. For instance, in evaluating Eq. (22) for $\Omega_{\mathrm{B}}^{d,2}$, we assume that $|\mu|\gg\varepsilon_{\max}$, where $\varepsilon_{\max}$ denotes the particle bandwidth, and that $|\mu|\gg|q_{\ell}|$ over the relevant bosonic energy scales   [^2] For the three-dimensional continuum model in the BEC limit, the condition $|\mu|\gg\Lambda^{2}/(2m)$ guarantees that the characteristic pair size $\xi_{\mathrm{pair}}$ is much smaller than the disorder correlation length [20]. Since $\xi_{\mathrm{pair}}\propto a_{s}$ and $|\mu|\simeq 1/(2ma_{s}^{2})$ in this regime, the required separation of length scales is automatically satisfied. An analogous argument applies to the lattice model, where in the strong-coupling limit $U/t\gg 1$ one has $|\mu|\simeq U/2$ and $\varepsilon_{\max}=12t$, while the pair size scales as $\xi_{\mathrm{pair}}\propto a\,t/U$. Under these conditions, the composite bosons respond to the disorder potential as effectively point-like particles . Under these conditions, we may approximate $\sum_{p}\mathcal{G}_{q+p}\mathcal{G}_{k+q+p}\mathcal{G}_{-p}\to\sum_{p}\mathcal{G}_{p}^{2}\mathcal{G}_{-p},$ and use Eq. (33), which leads to $\Omega_{\mathrm{B}}^{d,2}\to-\frac{2\kappa}{\beta N^{2}}\sum_{qk,k_{n}=0}\frac{\Gamma_{q}\Gamma_{q+k}}{Z^{2}}.$ Thus, we identify
@@
Thus, we conclude that while the leading-order self-energy diagrams associated with our bosonic contributions $\Omega_{\mathrm{B}}^{d,2}$ and $\Omega_{\mathrm{B}}^{d,1a}$ are properly included in Ref. [20], the diagram corresponding to the remaining bosonic term $\Omega_{\mathrm{B}}^{d,1b}$ is absent from their analysis. However, as discussed in Ref. [20], one finds that $\Sigma_{q}^{\mathrm{B}d,1a}\ll\Sigma_{q}^{\mathrm{B}d,2}$ in the BEC limit, and the same hierarchy holds for the remaining contribution $\Sigma_{q}^{\mathrm{B}d,1b}.$ Therefore, the leading-order effects of weak disorder near $T_{c}$ can be incorporated into the number equation through the self-energy as
@@ V Critical temperature $T_{c}$
It turns out that the effect of the disorder potential on $T_{c}$ differs markedly between a weakly-interacting Fermi gas in the BCS limit and a weakly-interacting Bose gas of tightly-bound fermion pairs in the BEC limit [19, 20]. While the self-energy expressions in Eqs. (24) and (26) may appear superficially similar, a key difference emerges: in the fermionic case, the chemical potential $\mu$ is large and positive at $T_{c}$, effectively shielding the system from disorder. As a result, $T_{c}$ is primarily determined by the saddle-point condition Eq. (12), and Anderson’s theorem protects the transition temperature from weak disorder at the level of the saddle-point and Gaussian fluctuation analysis. In other words, the presence of a large Fermi surface restricts scattering processes, ensuring that weak disorder has only a minimal effect. In contrast, for tightly-bound fermion pairs in the BEC limit, the effective bosonic chemical potential $\bar{\mu}_{M}\to 0^{+}$ at $T_{c}$, leaving them fully exposed to disorder. Since $T_{c}$ is essentially set by the number equation Eq. (37), rather than by Eq. (12) which fixes $\mu\to-\varepsilon_{b}/2$, weak disorder directly affects phase coherence through the emergence of an incoherent component in the pair propagator, characterized by the $\sqrt{-iq_{\ell}}$ dependence of the bosonic self-energy $\Sigma_{q}^{M}$. This distinction highlights the fundamental role of quantum statistics in determining the robustness of phase coherence against the disorder potential.
@@
On the other hand, in the BEC limit, direct evaluation of Eq. (34) gives $n_{\mathrm{B}}^{0}=2[m_{M}T_{c}/(2\pi)]^{3/2}\zeta(3/2),$ where $\zeta(3/2)=2.612$ is the Riemann-zeta function, and the leading-order disorder correction is given by Eq. (35). The latter can be calculated either by taking the derivative of Eq. (26) with respect to $\mu_{M}$, or equivalently by performing the momentum sum over $\mathbf{q}$ via direct integration, using $\int_{0}^{\infty}\frac{\sqrt{\varepsilon}\,d\varepsilon}{(iq_{\ell}-\varepsilon)^{2}}=\frac{\pi}{2\sqrt{-iq_{\ell}}}$ for $q_{\ell}\neq 0$. This yields $n_{\mathrm{B}}^{d}=\frac{\kappa_{M}m_{M}^{3}}{2\pi^{2}}T_{c},$ where we set $\bar{\mu}_{M}\to 0$ and employ $\sum_{q_{\ell}\neq 0}e^{iq_{\ell}0^{+}}=-1.$ The latter identity follows from $\frac{1}{\beta}\sum_{q_{\ell}}e^{iq_{\ell}0^{+}}=\frac{1}{2\pi i}\oint e^{z0^{+}}n_{\mathrm{B}}(z)dz=0,$ since the integrand vanishes along the circular periphery when the contour encircling the poles is extended to infinity. Note that the $q_{\ell}=0$ frequency does not contribute to the disorder correction, and the arising $1/\sqrt{-iq_{\ell}}$ factor is exactly canceled by the $\sqrt{-iq_{\ell}}$ dependence of the bosonic self-energy. See Appendix A for further discussion. As a result, we find
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: Self-consistent solutions for (a) the critical temperature $T_{c}$, (b) its disorder suppression $\delta T_{c}=T_{c,0}-T_{c}$ with $T_{c,0}$ for the clean case, (c) the disorder correction $n_{\mathrm{F}}^{d}+n_{\mathrm{B}}^{d}$ to the number equation, and (d) the chemical potential $\mu$. The inset in (d) shows the corresponding disorder correction $\delta\mu=\mu_{0}-\mu$, with $\mu_{0}$ for the clean case. All quantities are plotted as functions of the onsite interaction strength $U$ for $\kappa/t=\{0,0.02,0.05,0.1\}$ at quarter filling ($n=0.5$).
In the BCS-BEC crossover problem, one must solve the saddle-point condition in Eq. (12) together with the number equation in Eq. (27) self-consistently in order to determine $T_{c}$ and $\mu$ as functions of $U$. Even in the clean case, this procedure is already nontrivial in the crossover regime [23]; the inclusion of disorder renders the problem substantially more involved due to the complicated structure of the corrections appearing in Eqs. (20)-(22). In Ref. [20], a comprehensive numerical analysis of the resulting leading-order self-energy diagrams was carried out across the BCS-BEC crossover under various approximations, together with analytic estimates in the extreme BCS and BEC limits. It was further shown that incorporating disorder effects at the level of Eqs. (30) and (37) provides a qualitatively reliable description in the intermediate regime [19], although quantitative predictions depend sensitively on the specific procedure used to treat cutoff-related issues [20].
As an illustration, in Fig. 2 we consider a cubic lattice with dispersion
$\varepsilon_{\mathbf{k}}=2t\left[3-\cos(k_{x}a)-\cos(k_{y}a)-\cos(k_{z}a)\right],$ (39)
where $a$ is the lattice spacing and $t>0$ is the nearest-neighbor hopping amplitude. We present the self-consistent solutions for $T_{c}$, obtained by solving Eq. (12) together with Eqs. (30) and (37) at quarter filling. Furthermore, we approximate the effective dispersion of fermion pairs as $\varepsilon_{\mathbf{q}}^{M}=\frac{4t^{2}}{U}\left[3-\cos(q_{x}a)-\cos(q_{y}a)-\cos(q_{z}a)\right],$ which corresponds to an effective nearest-neighbor hopping $t_{M}=2t^{2}/U$ for the pairs (or, equivalently, an effective mass $m_{M}=U/(4t^{2}a^{2})$) and ensures that the interpolation scheme reproduces the correct BEC limit. Thus, while this numerical implementation captures both the BCS and BEC limits by construction, it provides only a qualitative description in the crossover regime. Consistent with our analysis above, the figure shows that the superfluid phase is most robust against disorder in the BCS and crossover regimes, whereas disorder significantly suppresses $T_{c}$ in the BEC regime. Although this trend is not clearly visible in the figures, we also observe a slight enhancement of $T_{c}$ in the weak-coupling BCS limit. For example, at $U=0.5t$, we obtain $T_{c}\approx 0.000935t$ in the clean case, while $T_{c}/t\approx\{0.000946,0.000950,0.000956\}$ for $\kappa/t=\{0.02,0.05,0.1\}$, respectively. However, we note that our numerics do not converge with high precision at such small values of $U/t$. We further find that the chemical potential $\mu$ is reduced relative to its clean-limit value $\mu_{0}$ in the BCS regime, consistent with the discussion in Sec. IV.1. In contrast, in the BEC regime $\mu$ is primarily determined by Eq. (12), rather than by Eq. (37), and therefore remains essentially unchanged from its clean-limit value.
Finally, we emphasize that while this Gaussian fluctuation framework provides a systematically controlled description near $T_{c}$ and recovers the correct BCS and BEC limits, it is not intended for a quantitatively exact treatment of the intermediate crossover. Since Gaussian theories often falter outside their controlled regimes, we employ this approximation to provide a consistent finite-temperature interpolation anchored by analytical limits, rather than a full self-consistent description of spectral properties across the entire crossover.